OpenStudy (anonymous):

divide radical 75 over 3?

zepdrix (zepdrix):

They both under the radical or no?

zepdrix (zepdrix):

\[\Large\rm \frac{\sqrt{75}}{3}\]Or\[\Large\rm \sqrt{\frac{75}{3}}\]

OpenStudy (anonymous):

Both under the radical

OpenStudy (anonymous):